OpenStudy (anonymous):

Can someone please help me with this factor the trinomial 21s^2-10s-24

OpenStudy (agreene):

(3s-4)(7s+6)

hero (hero):

Yeah, but you didn't post any of the steps

hero (hero):

vmart, would you like steps?

hero (hero):

I take that as a no

OpenStudy (anonymous):

i figured it out thank you so much!!

hero (hero):

The thing I worry about is you only figured it out by getting the answer. I'm not really convinced that you understood how to get there, but that's just me.

OpenStudy (anonymous):

(3s-4)(7s+6) = 21s^2 +18s-28s-24 = 21s^2-10s-24 = )

hero (hero):

Yes, that's only the "check" to make sure it's right.

hero (hero):

But all the steps are there. It looks to me like you were having trouble with factoring by grouping, and you needed the answer in order to figure out the steps.

OpenStudy (anonymous):

i probably will for the next problem any suggestions

hero (hero):

My suggestion is for you to learn how to factor by grouping. The first step in factoring by grouping is to split the middle term. There are several ways to approach it. One way is to figure out a "d" variable. (Yes, a "d" variable exists in addition to the a,b, and c variables.) There is a systematic approach that would allow to find the d value and properly split the mid-term. The second way to do it is to multiply a*c and continue to divide by factors of ac until you find two numbers that add to get b and multiply to get ac.

hero (hero):

When dividing by factors of ac, you normally divide ac by 1, then 2, then 3 .... and any numbers that are factors of ac until you find the right combination

hero (hero):

Or you can use agreene's intuitive method, which is difficult to explain

hero (hero):

But it is the fastest

OpenStudy (anonymous):

The first explanation is good thank you both hero & agreene i appreciate it !

hero (hero):

I like factoring by grouping because it's systematic and allows me to show my steps algebraically.

OpenStudy (anonymous):

thats nice to know